OpenStudy (anonymous):

The region of the solar system between Mars and Jupiter contains many asteroids that orbit the Sun. Consider an asteroid in a circular orbit of radius 4.0 1011 m. Find the period of the orbit

OpenStudy (anonymous):

4.0e11

OpenStudy (anonymous):

The question is presented as follows: No further info is given The region of the solar system between Mars and Jupiter contains many asteroids that orbit the Sun. Consider an asteroid in a circular orbit of radius 4.0e11 m. Find the period of the orbit

OpenStudy (anonymous):

@roadjester any ideas?

OpenStudy (anonymous):

Yeah..I'm working on it...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

use this equation which assumes the mass orbiting is negligible compared to the sun..which it is: Vo = sq rt [ GM /r] G = gravitational acceleration of sun = 274 meters/second M = mass of sun = 1.9891×10^30 kg Vo = velocity of orbiting asteroid r = radius of orbit = 4.0 * 10^11 meters Velocity of orbit = sq rt [(( 274 )(1.9891*10^30)) / (4*10^11)] Velocity of asteroid = 3.69125 * 10^10 m/s Now you need the circumference of that orbit and then the period is the time it takes for that asteroid to circle the sun once: Circum = 2* pi * r Circum = 2 *3.14 * 4*10^11 circumference of asteroid orbit = 2.51327 * 10^12 meters And last, divide distance of orbit by velocity of orbit to get period: (2.51327 * 10^12 m) / (3.69125 * 10^10 m/s) period = 68 seconds That is one fast asteroid

OpenStudy (anonymous):

gravitational acceleration of sun = 274 m/s^2....sorry

OpenStudy (mrnood):

According to Kepler's Third Law, the orbital period T\, (in seconds) of two bodies orbiting each other in a circular or elliptic orbit is: T = 2\pi\sqrt{a^3/\mu} where: a\, is the orbit's semi-major axis, in kilometers \mu = GM \, is the standard gravitational parameter, typically in km^3/s^2 G \, is the gravitational constant, M \, is the mass of the more massive body. For all ellipses with a given semi-major axis the orbital period is the same, regardless of eccentricity.

OpenStudy (mrnood):

http://en.wikipedia.org/wiki/Orbital_period